Write the polynomial x^2+4x-3 as a linear combination of 2x^2-3x and x^2-2x+5 and x+3


1
Expert's answer
2021-08-04T15:28:07-0400

Let


"x^2+4x-3=a(2x^2-3x)+b(x^2-2x+5)+c(x+3)"


"x^2+4x-3=(2a+b)x^2+(-3a-2b+c)x+(5b+3c)"

"x^2:2a+b=1"

"x^1:-3a-2b+c=4"

"x^0:5b+3c=-3"

"a=\dfrac{1}{2}-\dfrac{1}{2}b"

"c=-1-\dfrac{5}{3}b"

"-\dfrac{3}{2}+\dfrac{3}{2}b-2b-1-\dfrac{5}{3}b=4"



"-\dfrac{13}{6}b=\dfrac{13}{2}"

"b=-3"


"a=2"


"c=4"


"x^2+4x-3=2(2x^2-3x)-3(x^2-2x+5)+4(x+3)"